1. A weak acid (benzoic acid) is titrated with a strong base such as sodium hydroxide. Determine the pH at the half equivalence point of the titration. The Ka of the weak acid is 6.3 x 10-5. 2.Calculate the pH of a buffer made from 0.6 M HNO2 and 0.5 M NaNO2. Ka = 4.5 x 10-5 pKa = 4.3

Answer with one digit after the decimal place (e.g. 2.1) Hint: HNO2 is a weak acid. NaNO2 forms the conjugate base NO2-

3. Calculate the equilibrium constant at 25 °C for the weak base (CH3)2NH. Its weak acid is (CH3)2NH2+ , Ka = 1.7 x 10-11 Please use decimal notation and use two digits
Answer #1

1)

at half equivalent point

PH = Pka

Ka = 6.3 x10^-5

-log(Ka) = -log( 6. 3 x10^-5)

PKa = 4.20

Hence, at half equivalent point PH = 4.20

2)

Concentration of HNO2 = 0.6M

Concentration of NaNO2 = 0.5M

Ka = 4.5 x10^-5

-log(Ka) = -log( 4.5 x10^-5)

Pka = 4.35

for acidic buffer

PH = Pka + log[salt]/[acid]

PH = 4.35 + log( 0.5 / 0.6)

PH = 4.27